OpenStudy (anonymous):

Solve for x. 11x + 4 37

OpenStudy (anonymous):

x=3

OpenStudy (arbershabani97):

11x=37-4 x=(37-4)/11 and solve

OpenStudy (anonymous):

thx @catjean